### 05 Jan 2024 v1

# Modeling of Proton Interaction with Organic Polymers: Implications for Cancer Therapy and Beyond

Abstract
This comprehensive study delves into the intricate interplay between protons and organic polymers, offering insights into proton therapy in cancer treatment. Focusing on the influence of the spatial electron density distribution on stopping power estimates, we employed time-dependent density functional theory (TDDFT), coupled with the Penn method. Surprisingly, the assumption of electron density homogeneity in polymers is fundamentally flawed, resulting in an overestimation of stopping power values at energies below 2 MeV, approximately. Moreover, Bragg’s rule application in specific compounds exhibited significant deviations from experimental data in the Bragg peak region, challenging established norms.
preprint: APS/123-QED
INTRODUCTION
In the last two decades, clinical therapy using proton beams for the treatment of cancerous tumors has experienced steady growth [1, 2, 3, 4]. Although this form of radiotherapy already counts with highly developed technology, it still retains significant challenges in terms of physical and clinical aspects [5, 6, 7]. One of these challenges is the precise accounting of relative biological effectiveness (RBE), which is the ratio between the doses required by two types of radiation to cause the same biological effect. This factor, measurable through Linear energy transfer (LET) or microdosimetry, depends on how the energy is deposited on a micrometric scale [8].
In proton therapy, RBE is traditionally defined by a constant value of 1.1 (relative to X-ray dose) for all points along the beam path and all stopping points [9, 10]. However, a comprehensive review of the available experimental data in the literature [11] reveals that, despite a lack of experimental standardization and, large uncertainties, there is evidence that RBE values vary considerably and can exceed 1.1 at the end of the beam range. These differences have clinical implications [12, 13], therefore, it is of great importance to reduce experimental uncertainties to accurately describe the effects of proton beams on tissues.
The energy of the proton transferred to the biological tissue is directly related to its velocity. As the proton slows down, the amount of energy transferred to the tissue increases, resulting in maximum dose deposition at a specific depth. This region around the peak of maximum dose deposition is known as the Bragg peak [14, 15, 16, 17]. It is the region of greatest interest in proton beam radiotherapy applications, and its precise positioning is crucial during the definition of the irradiation plan. This particular profile of proton beam energy deposition presents significant clinical advantages, especially for pediatric patients, by allowing optimal dose delivery to tumor tissue and by minimizing dose to organs at risk in surrounding areas, thus reducing the chances of future complications and induction of secondary tumors [18, 19, 20].
On the other hand, the high and relatively narrow dose peak makes both quality control in dose monitoring, and precise patient positioning, even more crucial, with the risk of damaging health tissues with high radiation dose. Therefore, in-depth investigations of the uncertainties in the range and stopping power values are essential for a more accurate dose distribution in patients [21, 22].
Radiation transport simulation codes such as FLUKA, MCNP and TOPAS are widely used for dose estimates in radiation therapy planning [23, 24, 25]. The simulation result enables dose calculations in the planning process, and its quality reflects the quality of fundamental data tables, such as stopping power, scattering cross sections, ionization probabilities, energy loss, and lateral straggling, among others.
The accurate knowledge of electronic stopping power, or electronic stopping cross-section (SCS), is not just essential in proton therapy. It is also important in many fields of science and technological applications, such as outer space exploration (space weathering), nanotechnology (ion beam modeling), material modifications, and nuclear fusion research (plasma-wall interaction) [26, 27, 28, 29, 30]. However, as depicted above, its most critical application lies in dosimetry for cancer treatment using ions, given the increasing global use of protons and heavier ions in radiation therapy and the risks involved [31, 32]. Therefore, SCS is a fundamental quantity that requires a detailed understanding of energy-loss processes.
In the realm of theoretical investigations into ion-matter interactions, it is customary to employ simplified models, exemplified by the homogeneous free electron gas (FEG) model, for the representation of valence electrons within materials. This pragmatic approach facilitates straightforward predictions of stopping power and yields results that closely align with experimental data [33, 34, 35, 36, 37, 38, 39, 40, 41].
Although the FEG model is reliable for materials with simple electronic structures, its effectiveness diminishes when dealing with materials characterized by complex electronic excitations. Here, we demonstrate that these materials can still be treated as a collection of FEG with high accuracy, ensuring simplicity and avoiding time-consuming full atomistic ab initio calculations.
For this purpose, we utilized stopping power for a FEG with different densities or plasmon frequencies from the calculations of the time-dependent density functional theory (TDDFT) [34, 42, 43, 39]. The results were averaged according to the Penn method [44].
In this framework, knowledge of the energy-loss function (ELF) of materials is essential. The Penn approach [45] introduced an algorithm to determine the electron inelastic mean free paths (IMFP) by utilizing a model dielectric function derived from the experimental ELF specific to the material under investigation. The same model has been applied to estimate the electron stopping power in various materials [46] and has been extended to calculate the non-linear stopping power of ions [44]. This extension involves using the ELF to appropriately weight contributions from different electron gas components within a statistical ensemble that characterizes the material of interest. The ELF functions at the optical limit can be found for different materials elsewhere [47].
For compounds such as hydrocarbons, Bragg’s rule has been used to calculate the stopping values of their constituent elements [14]. However, it can cause significant differences in stopping power in the Bragg peak region compared to the experimental data [48].
By applying the proposed formalism, we examined the validity of Bragg’s rule in the context of organic polymers. Specifically, we examined the cases of polyethylene (PE), polystyrene (PS), poly(2-vinylpyridine) (P2VP), polyacetylene (PA), poly(methyl methacrylate) (PMMA) and polyimide (PI). The study of these polymers is important because virtually all phantoms used for dose verification and quality assurance in proton therapy treatments are manufactured with polymers such as PMMA. Furthermore, some components that make up the proton accelerators are constructed with PE or PS [49, 50, 51, 52, 53, 54].
METHODS
TDDFT is a highly effective ab initio tool for describing electronic stopping power in spherical jelliums. The jellium model assumes a positive background (representing the ion cores) that provides a charge balancing for the electron gas. The advantage of such representation as compared to fully atomistic models is the computational efficiency. The TDDFT in a FEG has been shown to provide accurate results for near-free-electron systems whereas using an atomistic representation requires careful consideration of trajectories to calculate the random stopping power [55, 56].
The positive background density of the jellium with radius $R_{\rm cl}$ is defined by $n_{0}^{+}({\bf r})=n_{0}^{+}\Theta(r-R_{\rm cl})$, where $\Theta(x)$ denotes the Heaviside function. The electronic density of the cluster is determined by the Wigner-Seitz radius, $r_{s}$ ($4\pi r_{s}^{3}/3=1/n_{0}$). The total number of electrons in the cluster, $N_{e}$, is then given by $N_{e}=(R_{cl}/r_{s})^{3}$.
Although there have been minor refinements in terms of accuracy, the approach adopted in this work reflects the methodology used in [34, 57, 42, 39], and as such it will be briefly explained in this section. In this approach, the time evolution of electronic density incorporates, in a non-perturbative manner, the complete dynamic interaction between an external field and the medium. This computational framework has been used to analyze various issues in condensed matter systems, such as dynamic charge screening in metallic media [58], energy loss of atomic particles in matter [33, 34, 39], as well as many-body effects associated with hole screening in photoemission [57].
A static density functional theory (DFT) calculation is performed to obtain the system’s ground state. The time evolution of the complete electronic density, $n({\bf r},t)$, in response to an external field (in this case, a proton), is conducted within the framework of TDDFT in the Kohn-Sham regime (KS-TDDFT) (atomic units are used throughout, unless specified otherwise):
$i\frac{\partial\psi_{j}({\bf r},t)}{\partial t}~{}=~{}\left\{T+V_{\rm{eff}}([n],{\bf r},t)\right\}\psi_{j}({\bf r},t)~{}\mathrm{,}$ (1)
where $\psi_{j}({\bf r},t)$ are the Kohn-Sham orbitals and $T$ is the kinetic energy operator. The Kohn-Sham effective potential, $V_{\text{eff}}([n],{\bf r},t)$, is a function of the electronic density of the system: $n({\bf r},t)=\sum_{j\in occ.}{\left|\psi_{j}({\bf r},t)\right|^{2}}$. $V_{\text{eff}}$ is obtained as the sum of the external potential $V_{\text{ext}}^{+}$, the Hartree potential $V_{\text{H}}$, and the exchange-correlation potential $V_{\text{xc}}$: $V_{\rm{eff}}=V_{\rm{ext}}^{+}+V_{\rm{H}}+V_{\rm{xc}}$. $V_{\text{ext}}^{+}$ is the potential created by the proton. $V_{\text{xc}}({\bf r},t)$ is treated within a standard adiabatic local density approximation (ALDA) approach. The numerical procedure is very similar to that employed in Refs. [58, 33, 34, 59], where additional details can be found.
The energy loss is calculated by integrating the time-dependent induced force over the proton:
[…]
where $v$ is the (constant) velocity at which the proton traverses the jellium. Once the induced force on the proton is calculated, the average or effective stopping power is computed as the energy loss per unit path length, i.e.
$\left(\frac{dE}{dz}\right)_{\text{TDDFT}}=\frac{E_{\rm{loss}}}{2R_{cl}}.$ (3)
Recently, an alternative non-linear method has been introduced to characterize the stopping power of light and heavy ions in materials [44]. This method incorporates the influence of non-free electron distributions within a theoretical model for stopping power calculations, such as TDDFT. For a low energy proton ($v<v_{F}$, where $v_{F}$ is Fermi velocity), the Penn approach has been used recently in the transport cross section (TCS) [40]. This approach considers the combination of electron-gas responses characterized by nonuniform densities, similar to the approach outlined in the Penn method [45].
To achieve this goal, each free electron density is analyzed based on the material’s ELF at the optical limit, as follows:
$g(\omega_{p})=\frac{2}{\pi\omega_{p}}\text{ELF}(\omega_{p}).$ (4)
The stopping power depends on the plasmon frequency $\omega_{p}$, a value determined by the individual electron gas contributions obtained from $r_{s}$. $\omega_{p}=\sqrt{3}r_{s}^{-3/2}$. Therefore, the stopping power is now calculated as follows:
[…]
In the above equation, the term $\left(dE/dz\right)_{\text{TDDFT}}$ is calculated in the TDDFT framework using Equation (3). Because of that, we named this approach TDDFT-Penn.
RESULTS AND DISCUSSION
TDDFT-Penn calculations were performed according to Equations (3) to (5), and the electronic SCS results for PE, PS, P2VP, PA, PMMA, and PI to energetic protons are presented in Figure 2 to 7, respectively. The data used to calculate SCS with the TDDFT-Penn method are listed in Table 1 and the optical-ELF data for each polymer are shown in Figure 1.
Table 1: Data used in the TDDFT-Penn approach to calculate the electronic SCS of different polymers based on their monomers [60, 61, 62].
Polymer
Formula
ELF
range (eV)
Total/valence
electrons
$\rho$
[…]
We compare the results of our approach with ICRU49 [63], ICRU37 [64], SRIM-2013 [65], and TDDFT with the homogeneous assumption. This comparison shows the need to completely break down the assumption of spatial homogeneity of the valence electron density in complex materials, such as polymers. For example, the homogeneous assumption leads to an overestimation of the SCS values for proton energies below approximately 2000 keV, compared to the SRIM-2013 data.
Figure 1: Optical-ELF data for PE, PS, P2VP, PA, PMMA, and PI obtained from [60, 61, 62] and used to calculate electronic SCS with the TDDFT-Penn approach.
Our approach produces more realistic values and, at the same time, offers a physically sound approach to dealing with inhomogeneities.
Figure 2: Comparative electronic stopping cross-section (SCS) of protons within PE polymer. The blue short dash-dot line is the TDDFT results using a FEG unique ($r_{s}=1.75$ au). The cyan short dash line is the TDDFT-Penn approach result. Experimental data (uppercase letters) are within the Bragg peak region [66]. Semi-empirical models ICRU49 [63] and SRIM-2013 [65] are depicted.
We also included in the comparison the experimental data available at the IAEA database [66] (upper case letters), with which the TDDFT-Penn results show an excellent agreement, as well as with semi-empirical SRIM-2013 [65] results (red dashed line) using Bragg’s rule [14], as can be seen in Figures 2 and 3.
Figure 3: Illustration of electronic stopping cross-section (SCS) for protons in PS polymer. TDDFT results employing a FEG unique ($r_{s}=1.66$ au) are represented by the blue short dash-dot line. The cyan short dash line is the result TDDFT-Penn approach. Experimental data (uppercase letters) are concentrated in the Bragg peak area. Dielectric formalism results are indicated by the purple dash-dot line [62]. Semi-empirical models ICRU49 [63] and SRIM-2013 [65] are showcased.
Finally, SCS results from the dielectric formalism, particularly the Mermin-Energy-Loss-Function Generalized Oscillator Strength model (MELF-GOS) [62], are also included in the comparisons. This approach also considers inhomogeneities in the electron density of the material utilizing a similar experimental ELF and thus both methods will give similar mean excitation energies (I) (occurring in the Bethe formula for fast projectiles). However, it underestimates the SCS at low energies by a significant amount. Several studies use this approach to calculate SCS in biological media [62, 50, 67]. Because this theoretical model is linear, it loses accuracy for ion energies in the Bragg peak region. In this energy range, the non-linear effects become significant. Even though we have not presented MELF-GOS results for PE, we expect similar behavior as observed in the others.
Notably, the TDDFT-Penn results for PS (refer to Figure 3) exhibit a better agreement with the experimental data compared to SRIM-2013. SRIM-2013 employs Bragg’s rule, resulting in an excitation energy (I) for PS of 65.5 eV [68]. However, electron energy loss spectroscopy (EELS) data from experiments [61] for the compound PS suggest a lower I value (59.3 eV). This observation explains the higher values obtained with our approach in the Bragg peak region.
SCS results for P2VP and PA (refer to Figures 4 and 5) exhibit good agreement with SRIM-2013.
Figure 4: Display of electronic stopping cross-section (SCS) in P2VP polymer for protons. The blue short dash-dot line represents TDDFT results utilizing a unique FEG ($r_{s}=1.66$ au). The cyan short dash line is the result TDDFT-Penn approach. The purple dash-dot line signifies results based on dielectric formalism [62]. Semi-empirical model SRIM-2013 [65] is included.
However, it is important to note that we did not find any experimental data for comparison in this context. Additionally, the dielectric approach, employing a similar
Figure 5: Depiction of electronic stopping cross-section (SCS) of protons interacting with PA polymer. TDDFT results using a FEG unique ($r_{s}=1.62$ au) are represented by the blue short dash-dot line. The cyan short dash line is the result TDDFT-Penn approach. Results based on dielectric formalism are represented by the purple dash-dot line [62]. Semi-empirical model SRIM-2013 [65] is presented.
ELF function as input [62], deviates significantly below the Bragg peak.
Interestingly, for PMMA and PI (see Figures 6 and 7) Bragg’s rule indicates 74 eV [63] and 79.6 eV [64], while the experimental ELF [62] points to significantly lower values of 66 eV and 68 eV, respectively. It is worth noting that the core and bond (CAB) correction on SRIM-2013 [65] is small but makes the deviation from our approach even higher pointing to a correction in the opposite direction.
Figure 6: Comparison of electronic stopping cross-section (SCS) for protons interacting with PMMA polymer. TDDFT results using a FEG unique ($r_{s}=1.74$ au) are represented by the short blue dashed line. The cyan short dashed line is the result of the TDDFT-Penn approach. Results based on the dielectric formalism are displayed on the purple dashed-dotted line [62]. Semi-empirical models ICRU49 [63] and SRIM-2013 [65] with Bragg rule (red dashed line) and CAB (dotted green line) correction are showcased.
[…]
In an analogous fashion, in TiN compounds there is a transfer of 1.51 electrons from titanium to nitrogen [40], and the transferred charges. The transfer of charges is expected to be more noticeable in the double bonds between carbon and oxygen in polymers like PMMA and PI. Therefore, our results suggest the possibility of charge transfer occurring from carbon to oxygen or nitrogen on these polymers. If this transfer is indeed occurring, employing Bragg’s rule will likely lead to a decreased accuracy in predicting the SCS.
Figure 7: Visualization of electronic stopping cross-section (SCS) for protons interacting with PI polymer. TDDFT results using a FEG unique ($r_{s}=1.61$ au) are represented by the blue short dash-dot line. The cyan short dash line is the result TDDFT-Penn approach. The purple dash-dot line represents results based on dielectric formalism [62]. Semi-empirical models ICRU37 [64] and SRIM-2013 [65] with Bragg rule (red dashed line) and CAB (dotted green line) correction are presented.
Indeed, the SRIM-2013 results (red dashed line) for the PMMA and PI compounds are numerically lower in the Bragg peak region compared to TDDFT-Penn predictions.
Theorists have been working for decades to develop methods that can model the physical processes responsible for electronic SCS. Although these methods have been successful, they have yet to be able to cover a wide range of energy using a single approach. Some models explain a limited energy range, while others achieve good agreement with experimental data by treating inner and valence electrons differently. In this work, we present a theoretical framework based on ab initio calculations of different FEGs and the Penn method. This framework achieves excellent agreement with experimental data and reference data tables across a wide range of energy, using a single approach for all electrons.
To showcase the efficiency of our approach, we conducted research on the stopping power of organic polymers for protons. In general, a single FEG model is not enough to accurately describe the stopping power at the Bragg peak. Our new method provides a theoretical framework to tackle electron density variations. We have demonstrated the effectiveness of our method by obtaining excellent agreement with the experimental data and reference data tables for polymers such as PE, PS, P2VP, PA, PMMA and PI (as shown in Figures 2 to 7). Our findings highlight the importance of considering the intricate electronic structures of polymers in the theoretical modeling of stopping power.
These findings emphasize the importance of considering the complex electronic structures of polymers in predicting stopping power accurately. The agreement between the TDDFT-Penn approach and the experimental data, which are indicated by uppercase letters in Figures 2 and 3, demonstrates the precision of the present theoretical framework. However, further validation is necessary, particularly for polymers with limited data.
Differences in predicted mean excitation energies and SCS magnitudes between the TDDFT-Penn approach and semi-empirical models such as SRIM-2013, particularly in the Bragg peak region, highlight the potential influence of molecular structure on these predictions. The distinct chemical compositions of PMMA and PI, which have varying electronegativities in their bonds, likely contribute to charge transfer effects not adequately accounted for by Bragg’s rule. The primary consequence of the breakdown of Bragg’s rule is a reduction in the mean excitation energy by approximately 10% and a decrease in the projectile range by about 3 mm for 200 MeV protons.
In conclusion, the results of this research challenge established assumptions and emphasize the need for precise modeling of materials with complex electronic structures. This study enhances our understanding of ion-polymer interactions and provides a solid foundation for future applications in proton therapy and other fields where accurate predictions of stopping power are essential.
ACKNOWLEDGMENTS
This work was partially supported by IPEN (project number 2020.06.IPEN.32) and CNPq (project number 406982/2021-0). The authors acknowledge FAPESP for supporting the computer cluster (process numbers 2012/04583-8 and 2020/04867-2). TFS and MHT acknowledge the financial support provided by CNPq-INCT-FNA (project number 464898/2014-5).